How many liters of pure water must be added to 6 liters of a 30% solution to get a 10% solution
Answer by stanbon(75887)

How many liters of pure water must be added to 6 liters of a 30% solution to get a 10% solution
----
Equation:
water + water = water
x + 0.70*6 = 0.90(x+6)
----
Multiply thru by 100 to get:
100x + 70*6 = 90x + 90*6
---
10x = 20*6
x = 3 liters (amt. of pure water needed)
===========
Cheers,
Stan H.
===========